RC meeting and Run Plan, Friday February 24th

From SBS wiki
Jump to: navigation, search

RC Report


  • Beam Delivery was reasonable from middle of yesterday Swing till start of Day today
  • Got 1.34 C in the last 24 hrs

Recent Issues

  • Slight problem with scaler GUI and DAQ due to sbsvme29. Resolved pretty quickly by rebooting the crate

Current Status

  • Machine is down for work in North Linac and Hall D Dipole

Run Plan

  • Continue running 45 uA on He3

Data Collection and Polarization Status

  • 41.044 C of uncorrected charge for this kinematic (48.399C of corrected charge which is 30% of the plan)
  • Last NMR 49.1%

Other Topics/Status Updates

  • Readiness target replacement
  • Moller polarimeter readiness
  • 6 Shifts still need filled during the remaining run period

Requested Accesses to the Hall